Find the RREF of the following equations:
3x + 4y - z = 7
x - 2y + 4z = -2
2x + 3y + 2z = 2

Answer by stanbon(75887)

Find the RREF of the following equations:
3x + 4y - z = 7
x - 2y + 4z = -2
2x + 3y + 2z = 2
-------
Using my TI-84 I get:
x = 2
y = 0
z = -1
============
Cheers,
Stan H.
============